OpenStudy (crashonce):

The three sides of a right angled triangle are of integral lengths which form an AP. Which of the following could be the length of one of the sides? 22, 58, 81, 91, 361 @mayankdevnani

OpenStudy (mayankdevnani):

 use pythagoras theorem

OpenStudy (crashonce):

ooooooo

OpenStudy (mayankdevnani):

\[\large \bf a^2+(a-d)^2=(a+d)^2\] \[\large \bf a^2+a^2+d^2-2ad=a^2+d^2+2ad\]

OpenStudy (mayankdevnani):

\[\large \bf a^2=4ad\] \[\large \bf a=4d\]

OpenStudy (crashonce):

im not sure how that helps tho

OpenStudy (mayankdevnani):

if we take a=22,then d=22/4=5.5

OpenStudy (mayankdevnani):

but we have to take only integral lengths So, 22 is not the answer

OpenStudy (mayankdevnani):

understood ??

OpenStudy (crashonce):

ahhhhhhhhhh yes

OpenStudy (mayankdevnani):

so,next option :- a=58,d=58/4=14.5

OpenStudy (crashonce):

nope

OpenStudy (mayankdevnani):

so this is also not our answer

OpenStudy (crashonce):

none of them work...

OpenStudy (mayankdevnani):

next,a=81 then d=81/4=20.25

OpenStudy (mayankdevnani):

not our answer again !!

OpenStudy (mayankdevnani):

then,a=91 So,d=91/4=22.75

OpenStudy (mayankdevnani):

not our answer

OpenStudy (mayankdevnani):

now last a=361 d=361/4=90.25

OpenStudy (crashonce):

361 doesnt work either?

OpenStudy (mayankdevnani):

this is also not our answer

OpenStudy (mayankdevnani):

i think no one satisifes

OpenStudy (mayankdevnani):

do you know the correct answer ??? @CrashOnce

OpenStudy (crashonce):

no i dont

OpenStudy (crashonce):

should we try again with a, a+d and a+2d

OpenStudy (mayankdevnani):

show your work !!

OpenStudy (mayankdevnani):

and i am thinking !! wait.

OpenStudy (crashonce):

no it wont work

OpenStudy (mayankdevnani):

i think nothing is the answer

OpenStudy (mayankdevnani):

@Directrix

OpenStudy (mayankdevnani):

@dan815 @triciaal

OpenStudy (mayankdevnani):

@jim_thompson5910 @ikram002p

OpenStudy (mayankdevnani):

@ganeshie8 @ganeshie8 plzzzz help...

OpenStudy (mayankdevnani):

@ganeshie8 @ganeshie8 @ganeshie8

OpenStudy (ikram002p):

well yes according to @mayankdevnani a=4d a+d=5d a-d=3d so , try all options ,

OpenStudy (ikram002p):

22,dont divide 3,5,4 58, dont divide 3,5,4 81,divide 3 , so it might be a-d 91, dont divide 3,4,5 361 dont divide 3,4,5,